Determine whether the following are propositions. Give the truth value of the propositions.

a) Asia is the largest of the Earth’s seven continents.

b) 4 + x = 5

c) 2 * 3 = 5

d) Answer this question

e) The human heart is shaped like an upside-down pear and is located slightly to the left of center inside the chest cavity


1
Expert's answer
2022-03-25T15:33:30-0400

Proposition is an expression that can be classified as truth or false. Expression has truth value 1 if it's true and 0 if false

a) Asia is indeed the largest of the Earth’s seven continents, so it is proposition with truth value 1

b) 4+x=5 cannot be determined as true or false, it depends on value of x. not a proposition

c) 2*3=5 is a proposition with truth value 0, cause 2*3=6

d) This sentence cannot be determined as true or false. not a proposition

e) The human heart is indeed shaped like an upside-down pear and is located slightly to the left of center inside the chest cavity, so it is proposition with truth value 1